OpenStudy (anonymous):

find the derivative f(x)=tan^(-1) (x/7)

OpenStudy (anonymous):

Do you know how to find the derivative of tan^(-1) ?

OpenStudy (anonymous):

1/x^2+1

OpenStudy (anonymous):

let's use the chain rule.

OpenStudy (anonymous):

\[y = \tan^{-1}(u)\]

OpenStudy (anonymous):

\[u = \frac{x}{7}\]

OpenStudy (anonymous):

we know that \[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}\]

OpenStudy (anonymous):

So \[\frac{dy}{dx}=\frac{1}{1+u^2}\frac{1}{7}\]

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{1}{1+(x/7)^2}\frac{1}{7}\]

OpenStudy (anonymous):

Do you have any question?

OpenStudy (anonymous):

its that whole thing right? 1/1+(x/7)^2*1/7???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome.